OpenStudy (anonymous):

What is the solution of |2x-1|=5 ?

OpenStudy (lgbasallote):

solve for x when 2x - 1 = 5 and 2x - 1 = -5

OpenStudy (anonymous):

so those bars dont really mean anything your still just solving for x?

OpenStudy (lgbasallote):

nope notice what i did

OpenStudy (lgbasallote):

those bars mean something

OpenStudy (lgbasallote):

you solve for \[\pm (2x -1) = 5\]

OpenStudy (anonymous):

ohh okay! thanks :) again.. lol:)

OpenStudy (lgbasallote):

welcome ^_^